HW Assignment 1 CLICK HERE TO REVIEW LEARNING OBJECTIVES QUESTON 8 Answer saved Marked out of 2.00 P ag question Determining Stockholders Equity and Assets Determine the following a. The stockholders equity of a corporation that has assets of $550,000 and liabilities of $306,000 s 0 b. The assets of a corporation that has liabilities of $250,000, common stock of $100,000, and retained earnings of $85,000. s 0 Save Answers Previous page Finish attempt MacBook Air 80 5 0
Answer #1

(a)

Assets = Liabilities + Stockholders' equity

550,000 = 306,000 + Stockholders' equity

Hence, Stockholders' equity = 550,000 - 306,000

= $244,000

(b)

Assets = Liabilities + Stockholders' equity

= 250,000 + (100,000 + 85,000)

= 250,000 + 185,000

= $435,000

Kindly give a positive rating if you are satisfied with the answer. Feel free to ask if you have any doubts. Thanks.